OpenStudy (anonymous):

2=3r+2+r solve each equation.

OpenStudy (anonymous):

2=3r+2+r=4r+2 4r=0 r=0

OpenStudy (anonymous):

2=3r+2+r subtract 2 from both sides 0=3r+r add the r's together 0=4r so r must = 0